OpenStudy (anonymous):

Sqrt [ (2^2 * 5 * ((x^3)^2) ] / sqrt 5^2 simplyfy please

OpenStudy (anonymous):

can you help me @Flop

OpenStudy (anonymous):

whats \[\sqrt{x ^{2}}\]?

OpenStudy (anonymous):

wait please

OpenStudy (anonymous):

\[\sqrt{2^{2} \times 5 \times (x ^{3})^{2}} \div \sqrt{5^{2}}\]

OpenStudy (anonymous):

this is the equation

OpenStudy (anonymous):

I know. Im trying to make a point using my problem. Do you know the answer to my question?

OpenStudy (anonymous):

yes i know square root of x to the power of 2 = x

OpenStudy (anonymous):

k. knowing that, the \[\sqrt{5^{2}}\] is just 5, 2^2=4, 4*5=20, (x^3)^2=x^6....so what does the final equation look like? There is one more simplification you can make to \[\sqrt{20}\]

OpenStudy (anonymous):

and also \[\sqrt{x ^{6}}\] can be simplified further. Final answer should be \[x^{3}2\sqrt{5}\div5\]

OpenStudy (anonymous):

thank you :)